# Efficient Defection: Overage-Proportional Rationing Attains the Cooperative Frontier

### v2

@@ 3 Main properties
With $C_{j}\in[0,M]$, for any fixed $C_{-j}$ the map $C_{j}\mapsto\pi_{j}(C_{j},C_{-j})$ is nondecreasing; thus $C_{j}^{\star}=M$ is a best reply independent of $C_{-j}$. (Appendix C.)
@@ Appendix A. Overage-power family and boundary continuity
$\widehat{v}_{j}^{\alpha}(C)=\begin{cases}v_{j},&X<I,\\[6.0pt] \dfrac{I\,v_{j}^{\alpha}}{\sum_{m:v_{m}>0}v_{m}^{\alpha}},&X>I,\\[10.0pt] \max\!\Big{\{}\,v_{j},\;\dfrac{I\,v_{j}^{\alpha}}{\sum_{m:v_{m}>0}v_{m}^{\alpha}}\,\Big{\}},&X=I\quad\text{(boundary tie-break)},\end{cases}$ (3)
@@
The tie-break ensures global monotonicity in own claim for every $\alpha>0$; for $\alpha=1$ the two branches coincide at $X=I$, so the rule is continuous without tie-break.
@@
with the boundary value at $X_{-j}+y=I_{-j}$ set to the maximum of the two (tie-break above). On $0,(I_{-j}-X_{-j})_{+}]$ we have $\pi_{j}^{\alpha}(y)=L_{j}+y$ which is strictly increasing. On $((I_{-j}-X_{-j})_{+},\;M-L_{j}]$ we have
@@
so $\pi_{j}^{\alpha}$ is nondecreasing there. By the tie-break at the boundary and the two regional conclusions, $\pi_{j}^{\alpha}(\cdot,C_{-j})$ is nondecreasing on $[0,M]$. In particular a best reply is attained at $C_{j}^{\star}=M$ for every $\alpha>0$.
@@
Continuity away from $X=I$ is immediate. At $X=I$ with positive $v$, approaching from $X<I$ gives $\widehat{v}_{j}^{\alpha}=v_{j}$ while from $X>I$ gives $\widehat{v}_{j}^{\alpha}=I\,v_{j}^{\alpha}/\sum_{\ell}v_{\ell}^{\alpha}$. Equality for all positive $v$ forces $\frac{v_{j}^{\alpha}}{\sum_{\ell}v_{\ell}^{\alpha}}=\frac{v_{j}}{\sum_{\ell}v_{\ell}}$, which holds iff $\alpha=1$; conversely, for $\alpha=1$ the branches coincide. The boundary tie-break does not affect this characterization. ∎

### v3

@@ 2 Mechanism (slack allocation)
Bounded actions.
When we study dominant-strategy behavior we restrict claims to $C_{j}\in[0,M]$ with a common bound $M>\max_{j}L_{j}$; this is the only place where boundedness is used (see Appendix C).
@@ 3 Main properties
With $C_{j}\in[0,M]$ and the linear rule ($\alpha=1$), for any fixed $C_{-j}$ the map $C_{j}\mapsto\pi_{j}(C_{j},C_{-j})$ is nondecreasing; thus $C_{j}^{\star}=M$ is a best reply independent of $C_{-j}$. (See Appendix C.)
@@ Appendix A. Overage-power family and boundary continuity
$\widehat{v}_{j}^{\alpha}(C)=\begin{cases}v_{j},&X\leq I,\\[6.0pt] \dfrac{I\,v_{j}^{\alpha}}{\sum_{m:v_{m}>0}v_{m}^{\alpha}},&X>I,\end{cases}$ (3)
@@
Boundary convention at $X=I$.
At the boundary we adopt the $X<I$ branch, i.e., $\widehat{v}_{j}^{\alpha}=v_{j}$. This preserves feasibility $\sum_{j:C_{j}>L_{j}}\widehat{v}_{j}^{\alpha}=I$ and budget balance. For $\alpha\neq 1$ the allocation generally differs from the $X>I$ branch, so a discontinuity remains; for $\alpha=1$ the two branches coincide and the rule is continuous.
@@
with the boundary value at $X_{-j}+y=I_{-j}$ set by the $X<I$ branch (boundary convention above).
On $0,(I_{-j}-X_{-j})_{+}]$ we have $\pi_{j}^{\alpha}(y)=L_{j}+y$ which is strictly increasing. On $((I_{-j}-X_{-j})_{+},\;M-L_{j}]$ we have
@@
so $\pi_{j}^{\alpha}$ is nondecreasing there. For $\alpha=1$, the branches coincide at the boundary, hence $\pi_{j}^{1}(\cdot,C_{-j})$ is nondecreasing on $[0,M]$. In particular a best reply is attained at $C_{j}^{\star}=M$ for $\alpha=1$.
@@
Continuity away from $X=I$ is immediate. At $X=I$ with positive $v$, approaching from $X<I$ gives $\widehat{v}_{j}^{\alpha}=v_{j}$ while from $X>I$ gives $\widehat{v}_{j}^{\alpha}=I\,v_{j}^{\alpha}/\sum_{\ell}v_{\ell}^{\alpha}$. Equality for all positive $v$ forces $\frac{v_{j}^{\alpha}}{\sum_{\ell}v_{\ell}^{\alpha}}=\frac{v_{j}}{\sum_{\ell}v_{\ell}}$, which holds iff $\alpha=1$; conversely, for $\alpha=1$ the branches coincide. The boundary convention does not affect this characterization. ∎